a man can type 21 words per minute.how long will it take him to type 3780 word?

Found 2 solutions by Theo, Shin123:
Answer by Theo(13342)

x = number of minutes.

21 * x = 3780

solve for x to get x = 3780 / 21 = 180.

the man would take 180 minutes to type 3780 words.


Answer by Shin123(626)

He can type 3780 words in (3780/21)=180 min./3 hrs.